OpenStudy (butterflyhope):

Help would be appreciated :) Solve and show work: y - 7 + 8y - 16 = -14 Thank you!

OpenStudy (anonymous):

add 16 to both sides

OpenStudy (anonymous):

then add 7 to both sides

OpenStudy (cj49):

9y-23=-14 9y=-14+23 9y=9 y=1

TheSmartOne (thesmartone):

Combine like terms 9y-23=-14 9y=9 y=1

OpenStudy (anonymous):

then add they y's

OpenStudy (anonymous):

and then divide